OpenStudy (anonymous):

-4t+1=3t+1-7t show step by step

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

\[-4t+1=3t-7t+1\] \[-4t+1=-4t+1\] \[-4t+4t=1-1\] \[0=0\] Hence no solution

OpenStudy (chaise):

-4t+1=3t+1-7t Gather like terms -4t+1=-4t+1 0/0 No solutions or infinite solutions.

OpenStudy (mimi_x3):

-4t + 1 = -4t + 1 No solution, since the t vanishes

OpenStudy (anonymous):

got it photo shop